The number of times the digit 3 will be written when listing the integers from 1 to 100 is


want full solution 


 


a)302   b)300   c)296    d)271

Grade:9

1 Answers

Aman Bansal
11 years ago

Dear Hrishikesh

Take all the number as  digits ( 01 to 99)
i fill the units place with 3 , i can fill the ten and hundreds place in 10 ways
10 * 10 * 1 = 100

Same way for hundreds = 10 * 1 * 10 = 100
thousands = 1* 10* 10 = 100

So totally 300 ways

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert